### v1

@@ 6 Unconditional Confidence Intervals
Suppose Assumptions 3.1 and 3.33.5 hold. Then, for any (random) $\hat{d}\in\{0,\ldots,T\}$ and $20<\alpha_{1},\alpha_{2}<1/$ 2,
@@
In analogy with the construction of the conditional CIs, to construct the hybrid CIs we begin by characterizing the conditional distributions of $\widehat{L}(\hat{d})$ and $\widehat{U}(\hat{d})$ but now adding an additional component to the conditioning events. More specifically, under Assumptions 3.2 and 3.6, by intersecting the events
@@
=$=\left\{\widehat{\mathcal{V}}_{L}^{-}\left(\widehat{Z}_{L}(d^{*},j_{L}^{*}),d^{*},j_{L}^{*},\gamma_{L}^{*}\right)\leq\sqrt{n}(\tilde{\ell}_{d^{*},j_{L}^{*}}+\ell_{d^{*},j_{L}^{*}}\hat{p})\leq\widehat{\mathcal{V}}_{L}^{+}\left(\widehat{Z}_{L}(d^{*},j_{L}^{*}),d^{*},j_{L}^{*},\gamma_{L}^{*}\right),\right.$
$\left.\widehat{\mathcal{V}}_{L}^{0}\left(\widehat{Z}_{L}(d^{*},j_{L}^{*}),d^{*},j_{L}^{*},\gamma_{L}^{*}\right)\geq 0\right\}$
@@
$\sim\left.\sqrt{n}(\tilde{\ell}_{d^{*},j_{L}^{*}}+\ell_{d^{*},j_{L}^{*}}\hat{p})\right|\left\{\widehat{\mathcal{V}}_{L}^{-}\left(\widehat{Z}_{L}(d^{*},j_{L}^{*}),d^{*},j_{L}^{*},\gamma_{L}^{*}\right)\leq\sqrt{n}(\tilde{\ell}_{d^{*},j_{L}^{*}}+\ell_{d^{*},j_{L}^{*}}\hat{p})\right.$
$\left.\leq\widehat{\mathcal{V}}_{L}^{+,H}\left(\widehat{Z}_{L}(d^{*},j_{L}^{*}),d^{*},j_{L}^{*},\gamma_{L}^{*},\sqrt{n}(\tilde{\ell}_{d^{*},j_{L}^{*}}+\ell_{d^{*},j_{L}^{*}}p)\right),\widehat{\mathcal{V}}_{L}^{0}\left(\widehat{Z}_{L}(d^{*},j_{L}^{*}),d^{*},j_{L}^{*},\gamma_{L}^{*}\right)\geq 0\right\},$
@@
$\widehat{\mathcal{V}}_{L}^{+,H}(z,d,j,\gamma,\mu)=\min\left\{\widehat{\mathcal{V}}_{L}^{+}(z,d,j,\gamma),\mu+\hat{c}_{1-\beta,L}\sqrt{\widehat{\Sigma}_{L,dJ_{L}+j}}\right\}.$
@@
$\sim\left.\sqrt{n}(\tilde{u}_{d^{*},j_{U}^{*}}+u_{d^{*},j_{U}^{*}}\hat{p})\right|\left\{\widehat{\mathcal{V}}_{U}^{-,H}\left(\widehat{Z}_{U}(d^{*},j_{U}^{*}),d^{*},j_{U}^{*},\gamma_{U}^{*},\sqrt{n}(\tilde{u}_{d^{*},j_{U}^{*}}+u_{d^{*},j_{U}^{*}}p)\right)\right.$
$\left.\leq\sqrt{n}(\tilde{u}_{d^{*},j_{U}^{*}}+u_{d^{*},j_{U}^{*}}\hat{p})\leq\widehat{\mathcal{V}}_{U}^{+}\left(\widehat{Z}_{U}(d^{*},j_{U}^{*}),d^{*},j_{U}^{*},\gamma_{U}^{*}\right),\widehat{\mathcal{V}}_{U}^{0}\left(\widehat{Z}_{U}(d^{*},j_{U}^{*}),d^{*},j_{U}^{*},\gamma_{U}^{*}\right)\geq 0\right\},$
@@
$\widehat{\mathcal{V}}_{U}^{-,H}(z,d,j,\gamma,\mu)=\max\left\{\widehat{\mathcal{V}}_{U}^{-}(z,d,j,\gamma),\mu-\hat{c}_{1-\beta,U}\sqrt{\widehat{\Sigma}_{U,dJ_{U}+j}}\right\}.$
Since the distribution of $\sqrt{n}(\tilde{\ell}_{d^{*},j_{L}^{*}}+\ell_{d^{*},j_{L}^{*}}\hat{p})$ can be approximated by a $\mathcal{N}(\sqrt{n}(\tilde{\ell}_{d^{*},j_{L}^{*}}+\ell_{d^{*},j_{L}^{*}}p),\ell_{d^{*},j_{L}^{*}}\Sigma\ell_{d^{*},j_{L}^{*}}^{\prime})$ distribution asymptotically and $\widehat{Z}_{L}(d^{*},j_{L}^{*})$ is asymptotically independent, we again work with the truncated normal distribution to compute a hybrid probabilistic lower bound for $\sqrt{n}L(\hat{d})$: for $10<\beta<\alpha<$ 1, define $\widehat{L}(d)_{\alpha}^{H}$ to solve
$F_{TN}\left(\sqrt{n}(\tilde{\ell}_{d,\hat{j}_{L}(d)}+\ell_{d,\hat{j}_{L}(d)}\hat{p});\mu,\ell_{d,\hat{j}_{L}(d)}\widehat{\Sigma}\ell_{d,\hat{j}_{L}(d)}^{\prime},\widehat{\mathcal{V}}_{L}^{-}\left(\widehat{Z}_{L}(d,\hat{j}_{L}(d)),d,\hat{j}_{L}(d),\hat{\gamma}_{L}(d)\right),\right.$
$\left.\widehat{\mathcal{V}}_{L}^{+,H}\left(\widehat{Z}_{L}(d,\hat{j}_{L}(d)),d,\hat{j}_{L}(d),\hat{\gamma}_{L}(d),\mu\right)\right)=\frac{1-\alpha}{1-\beta}.$
@@
$F_{TN}\left(\sqrt{n}(\tilde{u}_{d,\hat{j}_{U}(d)}+u_{d,\hat{j}_{U}(d)}\hat{p});\mu,u_{d,\hat{j}_{U}(d)}\widehat{\Sigma}u_{d,\hat{j}_{U}(d)}^{\prime},\widehat{\mathcal{V}}_{U}^{-,H}\left(\widehat{Z}_{U}(d,\hat{j}_{U}(d)),d,\hat{j}_{U}(d),\hat{\gamma}_{U}(d),\mu\right),\right.$
$\left.\widehat{\mathcal{V}}_{U}^{+}\left(\widehat{Z}_{U}(d,\hat{j}_{U}(d)),d,\hat{j}_{U}(d),\hat{\gamma}_{U}(d)\right)\right)=\frac{1-\alpha}{1-\beta}$
@@
with uniformly correct asymptotic coverage when $\hat{d}$ is selected from the data according to Assumptions 3.2 and 3.6.
@@ 7 Reality Check Power Comparison
In particular, we compare the power of a test implied by our hybrid CI (i.e., a test that rejects when the value of the parameter under the null hypothesis lies outside of the hybrid CI) to the power of the hybrid test of Andrews et al., (2023), which applies to a general class of moment-inequality models. When $d$ is chosen a priori and the parameter $p$ is equal to a vector of moments of underlying data, Assumption 3.1 implies that $L(d)\leq W(d)\leq U(d)$ can be written as a set of (unconditional) moment inequalities, a special case of the general framework of that paper. Of the many papers on inference for moment inequalities, we choose the test of Andrews et al., (2023) for comparison for two reasons: (i) it has been shown to be quite competitive in terms of power and (ii) it is also based upon a (different) inference method that is a hybrid between conditional and projection-based inference.
We compare the power of tests on the ATE of the treatment in the same setting of the Manski bounds example of Section 2, strengthening the mean independence assumption $E(Y(d)|Z)=E(Y(d))$ to full statistical independence $(Y(1),Y(0))\perp Z$ and using the sharp bounds on the ATE $W(1)=E(Y(1))-E(Y(0))$ derived by Balke and Pearl, (1997, 2011):
$L(1)=\max\left\{\begin{array}[]{c}p^{111}+p^{000}-1\\ p^{110}+p^{001}-1\\ p^{110}-p^{111}-p^{101}-p^{010}-p^{100}\\ p^{111}-p^{110}-p^{100}-p^{011}-p^{101}\\ -p^{011}-p^{101}\\ -p^{010}-p^{100}\\ p^{001}-p^{011}-p^{101}-p^{010}-p^{000}\\ p^{000}-p^{010}-p^{100}-p^{011}-p^{001}\end{array}\right\}$
@@
$U(1)=\min\left\{\begin{array}[]{c}1-p^{011}-p^{100}\\ 1-p^{010}-p^{101}\\ -p^{010}+p^{011}+p^{001}+p^{110}+p^{000}\\ -p^{011}+p^{111}+p^{001}+p^{010}+p^{000}\\ p^{111}+p^{001}\\ p^{110}+p^{000}\\ -p^{101}+p^{111}+p^{001}+p^{110}+p^{100}\\ -p^{100}+p^{110}+p^{000}+p^{111}+p^{101}\end{array}\right\}.$
@@ 8 Finite Sample Performance of Confidence Intervals
Moving now to a context in which the object of interest is selected from the data, we compare the finite sample performance of our conditional, projection and hybrid CIs again in the setting of the Manski bounds example of Section 2. In this case, we are interested in inference on the average potential outcome $W(\hat{d})$ for $W(d)=\mathbb{E}[Y(d)]$, where interest arises either in the average potential outcome for treatment ($\hat{d}=1$) or control ($\hat{d}=0$) depending upon which has the largest estimated lower bound: $\hat{d}=\operatorname*{argmax\;}_{d\in\{0,1\}}\widehat{L}(d)$. This form of $\hat{d}$ corresponds to case 1. of Proposition 3.1 and we use the corresponding result of the proposition to specify $\hat{\gamma}_{L}(\hat{d})$ and $\hat{\gamma}_{U}(d)$ in the construction of the conditional and hybrid CIs.
For the same DGPs as in Section 7, we compute the unconditional coverage frequencies of the conditional, projection and hybrid CIs as well as that of the conventional CI based upon the normal distribution. These coverage frequencies are reported in Table 1. Consistent with the asymptotic results of Theorems 5.16.2, the conditional, projection and hybrid CIs all have correct coverage for all DGPs and the modest sample size of $n=100$. Also consistent with Theorem 6.1, we note that the projection CI tends to be conservative with true coverage above the nominal level of 95%. Finally, we note that the conventional CI can substantially under-cover, consistent with the discussion in Section 2.1.
@@
This table reports unconditional coverage frequencies for the potential outcome selected by maximizing the estimated lower bound on the potential outcomes of either treatment or control, all evaluated at the nominal coverage level of 95%. Coverage frequencies are reported for conventional (“Conv”), conditional (“Cond”), projection (“Proj”) and hybrid (“Hyb”) confidence intervals for a sample size of $n=100$.
Next, we compare the length quantiles of the CIs with correct coverage for these same DGPs. Figure 2 plots the ratios of the $5^{th}$, $25^{th}$, $50^{th}$, $75^{th}$ and $95^{th}$ quantiles of the length of the conditional, projection and hybrid CIs relative to those same length quantiles of the projection CI. As can be seen from the figure, the conditional CI has the tendency to become very long, especially at high quantile levels, for certain DGPs, whereas the hybrid CI tends to perform the best overall by limiting the worst-case length performance of the conditional CI relative to the projection CI. Relative to projection, the hybrid CI enjoys length reductions of 20-30% for favorable DGPs while only showing length increases of 5-10% for unfavorable DGPs.

### v2

@@ 5.2 Unconditional Confidence Intervals
Suppose Assumptions 3.1 and 3.43.6 hold. Then, for any (random) $\hat{d}\in\{d^{0},\ldots,d^{K}\}$ and $20<\alpha_{1},\alpha_{2}<1/$ 2,
@@
In analogy with the construction of the conditional CIs, to construct the hybrid CIs we begin by characterizing the conditional distributions of $\widehat{L}(\hat{d})$ and $\widehat{U}(\hat{d})$ but now adding an additional component to the conditioning events. More specifically, under Assumptions 3.2 and 3.3, by intersecting the events
@@
=$=\left\{\widehat{\mathcal{V}}_{L}^{-}\left(\widehat{\mathcal{Z}}_{L}(d^{*},j_{L}^{*}),d^{*},j_{L}^{*},\gamma_{L}^{*}\right)\leq\sqrt{n}(\tilde{\ell}_{d^{*},j_{L}^{*}}+\ell_{d^{*},j_{L}^{*}}\hat{p})\leq\widehat{\mathcal{V}}_{L}^{+}\left(\widehat{\mathcal{Z}}_{L}(d^{*},j_{L}^{*}),d^{*},j_{L}^{*},\gamma_{L}^{*}\right),\right.$
$\left.\widehat{\mathcal{V}}_{L}^{0}\left(\widehat{\mathcal{Z}}_{L}(d^{*},j_{L}^{*}),d^{*},j_{L}^{*},\gamma_{L}^{*}\right)\geq 0\right\}$
@@
$\sim\left.\sqrt{n}(\tilde{\ell}_{d^{*},j_{L}^{*}}+\ell_{d^{*},j_{L}^{*}}\hat{p})\right|\left\{\widehat{\mathcal{V}}_{L}^{-}\left(\widehat{\mathcal{Z}}_{L}(d^{*},j_{L}^{*}),d^{*},j_{L}^{*},\gamma_{L}^{*}\right)\leq\sqrt{n}(\tilde{\ell}_{d^{*},j_{L}^{*}}+\ell_{d^{*},j_{L}^{*}}\hat{p})\right.$
$\left.\leq\widehat{\mathcal{V}}_{L}^{+,H}\left(\widehat{\mathcal{Z}}_{L}(d^{*},j_{L}^{*}),d^{*},j_{L}^{*},\gamma_{L}^{*},\sqrt{n}(\tilde{\ell}_{d^{*},j_{L}^{*}}+\ell_{d^{*},j_{L}^{*}}p)\right),\widehat{\mathcal{V}}_{L}^{0}\left(\widehat{\mathcal{Z}}_{L}(d^{*},j_{L}^{*}),d^{*},j_{L}^{*},\gamma_{L}^{*}\right)\geq 0\right\},$
@@
$\widehat{\mathcal{V}}_{L}^{+,H}(z,d,j,\gamma,\mu)\equiv\min\left\{\widehat{\mathcal{V}}_{L}^{+}(z,d,j,\gamma),\mu+\hat{c}_{1-\beta,L}\sqrt{\widehat{\Sigma}_{L,dJ_{L}+j}}\right\}.$
@@
$\sim\left.\sqrt{n}(\tilde{u}_{d^{*},j_{U}^{*}}+u_{d^{*},j_{U}^{*}}\hat{p})\right|\left\{\widehat{\mathcal{V}}_{U}^{-,H}\left(\widehat{\mathcal{Z}}_{U}(d^{*},j_{U}^{*}),d^{*},j_{U}^{*},\gamma_{U}^{*},\sqrt{n}(\tilde{u}_{d^{*},j_{U}^{*}}+u_{d^{*},j_{U}^{*}}p)\right)\right.$
$\left.\leq\sqrt{n}(\tilde{u}_{d^{*},j_{U}^{*}}+u_{d^{*},j_{U}^{*}}\hat{p})\leq\widehat{\mathcal{V}}_{U}^{+}\left(\widehat{\mathcal{Z}}_{U}(d^{*},j_{U}^{*}),d^{*},j_{U}^{*},\gamma_{U}^{*}\right),\widehat{\mathcal{V}}_{U}^{0}\left(\widehat{\mathcal{Z}}_{U}(d^{*},j_{U}^{*}),d^{*},j_{U}^{*},\gamma_{U}^{*}\right)\geq 0\right\},$
@@
$\widehat{\mathcal{V}}_{U}^{-,H}(z,d,j,\gamma,\mu)\equiv\max\left\{\widehat{\mathcal{V}}_{U}^{-}(z,d,j,\gamma),\mu-\hat{c}_{1-\beta,U}\sqrt{\widehat{\Sigma}_{U,dJ_{U}+j}}\right\}.$
Since the distribution of $\sqrt{n}(\tilde{\ell}_{d^{*},j_{L}^{*}}+\ell_{d^{*},j_{L}^{*}}\hat{p})$ can be approximated by $\mathcal{N}(\sqrt{n}(\tilde{\ell}_{d^{*},j_{L}^{*}}+\ell_{d^{*},j_{L}^{*}}p),\ell_{d^{*},j_{L}^{*}}\Sigma\ell_{d^{*},j_{L}^{*}}^{\prime})$ asymptotically and $\widehat{\mathcal{Z}}_{L}(d^{*},j_{L}^{*})$ is asymptotically independent, we again work with the truncated normal distribution to compute a hybrid probabilistic lower bound for $\sqrt{n}L(\hat{d})$: for $10<\beta<\alpha<$ 1, define $\widehat{L}(d)_{\alpha}^{H}$ to solve
$F_{TN}\left(\sqrt{n}(\tilde{\ell}_{d,\hat{j}_{L}(d)}+\ell_{d,\hat{j}_{L}(d)}\hat{p});\mu,\ell_{d,\hat{j}_{L}(d)}\widehat{\Sigma}\ell_{d,\hat{j}_{L}(d)}^{\prime}\right|\widehat{\mathcal{V}}_{L}^{-}\left(\widehat{\mathcal{Z}}_{L}(d,\hat{j}_{L}(d)),d,\hat{j}_{L}(d),\hat{\gamma}_{L}(d)\right),$
$\left.\widehat{\mathcal{V}}_{L}^{+,H}\left(\widehat{\mathcal{Z}}_{L}(d,\hat{j}_{L}(d)),d,\hat{j}_{L}(d),\hat{\gamma}_{L}(d),\mu\right)\right)=\frac{1-\alpha}{1-\beta}$
@@
$F_{TN}\left(\sqrt{n}(\tilde{u}_{d,\hat{j}_{U}(d)}+u_{d,\hat{j}_{U}(d)}\hat{p});\mu,u_{d,\hat{j}_{U}(d)}\widehat{\Sigma}u_{d,\hat{j}_{U}(d)}^{\prime}\right|\widehat{\mathcal{V}}_{U}^{-,H}\left(\widehat{\mathcal{Z}}_{U}(d,\hat{j}_{U}(d)),d,\hat{j}_{U}(d),\hat{\gamma}_{U}(d),\mu\right),$
$\left.\widehat{\mathcal{V}}_{U}^{+}\left(\widehat{\mathcal{Z}}_{U}(d,\hat{j}_{U}(d)),d,\hat{j}_{U}(d),\hat{\gamma}_{U}(d)\right)\right)=\frac{1-\alpha}{1-\beta}$
@@
with uniformly correct asymptotic coverage when $\hat{d}$ is selected from the data according to Assumptions 3.2 and 3.3.
@@ 6 Reality Check Power Comparison
In particular, we compare the power of the test implied by our hybrid CI (i.e., a test that rejects when the value of the parameter under the null hypothesis lies outside of the hybrid CI) to the power of the hybrid test of Andrews et al., (2023), which applies to a general class of moment-inequality models. When $d$ is chosen a priori and the parameter $p$ is equal to a vector of moments of underlying data, Assumption 3.1 implies that $L(d)\leq W(d)\leq U(d)$ can be written as a set of (unconditional) moment inequalities, a special case of the general framework of that paper. Of the many papers on inference for moment inequalities, we choose the test of Andrews et al., (2023) for comparison for two reasons: (i) it has been shown to be quite competitive in terms of power and (ii) it is also based upon a (different) inference method that is a hybrid between conditional and projection-based inference.
We compare the power of tests on the ATE in the same setting of the Manski bounds example of Section 2, strengthening the mean independence assumption $\mathbb{E}[Y(d)|Z]=\mathbb{E}[Y(d)]$ to full statistical independence $(Y(1),Y(0))\perp Z$ and using the sharp bounds on the ATE $W(1)-W(0)=\mathbb{E}[Y(1)]-\mathbb{E}[Y(0)]$ derived by Balke and Pearl, (1997, 2011):
$L=\max\left\{\begin{array}[]{c}p^{111}+p^{000}-1\\ p^{110}+p^{001}-1\\ p^{110}-p^{111}-p^{101}-p^{010}-p^{100}\\ p^{111}-p^{110}-p^{100}-p^{011}-p^{101}\\ -p^{011}-p^{101}\\ -p^{010}-p^{100}\\ p^{001}-p^{011}-p^{101}-p^{010}-p^{000}\\ p^{000}-p^{010}-p^{100}-p^{011}-p^{001}\end{array}\right\}$
@@
$U=\min\left\{\begin{array}[]{c}1-p^{011}-p^{100}\\ 1-p^{010}-p^{101}\\ -p^{010}+p^{011}+p^{001}+p^{110}+p^{000}\\ -p^{011}+p^{111}+p^{001}+p^{010}+p^{000}\\ p^{111}+p^{001}\\ p^{110}+p^{000}\\ -p^{101}+p^{111}+p^{001}+p^{110}+p^{100}\\ -p^{100}+p^{110}+p^{000}+p^{111}+p^{101}\end{array}\right\}.$
@@ 7 Finite Sample Performance of Confidence Intervals
Moving now to a context in which the object of interest is selected from the data, we compare the finite sample performance of our conditional, projection and hybrid CIs again in the setting of the Manski bounds example of Section 2. In this case, we are interested in inference on the average potential outcome $W(\hat{d})$ for $W(d)=\mathbb{E}[Y(d)]$, where interest arises either in the average potential outcome for treatment ($\hat{d}=1$) or control ($\hat{d}=0$) depending upon which has the largest estimated lower bound: $\hat{d}=\operatorname*{argmax\;}_{d\in\{0,1\}}\widehat{L}(d)$. This form of $\hat{d}$ corresponds to case 1. of Proposition 3.1 and we use the corresponding result of the proposition to specify $\hat{\gamma}_{L}(\hat{d})$ and $\hat{\gamma}_{U}(\hat{d})$ in the construction of the conditional and hybrid CIs. We report analogous simulation results for the dynamic treatment regime example in Appendix B with $\hat{p}$ generated from a multinomial, rather than normal, distribution.
For the same DGPs as in Section 6, we compute the unconditional coverage frequencies of the conditional, projection and hybrid CIs as well as that of the conventional CI based upon the normal distribution. These coverage frequencies are reported in Table 1. Consistent with the asymptotic results of Theorems 5.1, 5.2 and 5.3, the conditional, projection and hybrid CIs all have correct coverage for all DGPs and the modest sample size of $n=100$. Also consistent with Theorem 5.2, we note that the projection CI tends to be conservative with true coverage above the nominal level of 95%. Finally, we note that the conventional CI can substantially under-cover, consistent with the discussion in Section 2.1.
@@
This table reports unconditional coverage frequencies for the potential outcome selected by maximizing the estimated lower bound on the potential outcomes of either treatment or control, all evaluated at the nominal coverage level of 95%. Coverage frequencies are reported for conventional (“Conv”), conditional (“Cond”), projection (“Proj”) and hybrid (“Hyb”) CIs for a sample size of $n=100$.
Next, we compare the length quantiles of the CIs with correct coverage for these same DGPs. Figure 2 plots the ratios of the $5^{th}$, $25^{th}$, $50^{th}$, $75^{th}$ and $95^{th}$ quantiles of the length of the conditional, projection and hybrid CIs relative to those same length quantiles of the projection CI. As can be seen from the figure, the conditional CI has the tendency to become very long, especially at high quantile levels for certain DGPs, whereas the hybrid CI tends to perform the best overall by limiting the worst-case length performance of the conditional CI relative to the projection CI. Relative to projection, the hybrid CI enjoys length reductions of 20-30% for favorable DGPs while only showing length increases of 5-10% for unfavorable DGPs.